OpenStudy (anonymous):

In a shop there are 25 customers, 14 of whom will make a purchase. If three customers are selected, one at a time, at random, what is the probability that all will make a purchase? A) 0.1583 B) 0.1695 C) 0.2502 D) 0.2786

OpenStudy (anonymous):

A

OpenStudy (anonymous):

can you show me how? I have midterms tomorrow and need to know how

OpenStudy (anonymous):

probability for the first person chosen to make a purchase = 14/25 second person= 13/24 third person= 12/23

OpenStudy (anonymous):

so for all three to occur at the same time, u multiply all three together

OpenStudy (anonymous):

oh thats whats i missed divide by 3 thank oyu

OpenStudy (anonymous):

np